OpenStudy (anonymous):

Find f^-1 for the function f(x) = (1-41x)/x

OpenStudy (john_es):

You can find the path to the solution using this "trick" \[y=\frac{1-41x}{x}\Rightarrow x'=\frac{1-41y'}{y'}\] And solving for y'.

OpenStudy (jdoe0001):

\(\bf f(x) = \cfrac{1-41x}{x} \implies \color{red}{y} = \cfrac{1-41\color{blue}{x}}{\color{blue}{x}}\\\quad \\ f^{-1}(x) =\color{blue}{ x} = \cfrac{1-41\color{red}{y}}{\color{red}{y}} \implies x = \cfrac{1}{y}- \cfrac{41y}{y} \) you find the inverse function by "switching about" the variables

OpenStudy (anonymous):

would i multiply both sides by y to get rid of the denominator?

OpenStudy (jdoe0001):

keep in mind that \(\bf x = \cfrac{1}{y}- \cfrac{41\cancel{y}}{\cancel{y}} \implies x = \cfrac{1}{y} -41\)

OpenStudy (anonymous):

would it be y^-1 = 1/(x+ 41)

OpenStudy (jdoe0001):

yeap \(\bf x = \cfrac{1}{y}- \cfrac{41y}{y} \implies x = \cfrac{1}{y} -41 \implies y = \cfrac{1}{x+41}\)

OpenStudy (anonymous):

Awesome, thanks! I have one more like this if you would please check my answer. Find f^-1 for the function f(x) = (x-9)^3 + 3.

OpenStudy (anonymous):

I got \[\sqrt[3]{(x-3)} + 9\]